Circular No. 6913

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980519 A. Maury and D. Albanese, Observatoire de la Côte d’Azur (OCA);

M. Boer, Centre d’Etude Spatiale des Rayonnements, CNRS; and M. Feroci, Istituto Astrofisica Spaziale, CNR, report: “We obtained images of the field of GRB 980519 (IAUC 6909, 6910, 6912) on May 19 and 20 with the OCA 0.90-m Schmidt telescope. The optical transient reported by Jaunsen et al. (GCN 78) is clearly detected and fading on unfitered exposures taken on May 19.86256, 20.07652, and 20.96426 UT, at which time we measure magnitudes 19.06 ± 0.26, 19.81 ± 0.37, and 22.06 ± 0.76, respectively. The power-law slope decline is roughly 2.0, which, given the uncertainties, is consistent with the value reported by Djorgovski et al. (GCN 79). The images and the light curve are available at http://wwwrc.obs-

azur.fr/schmidt/980519.html.”

COMET C/1998 J1 (SOHO) Improved parabolic orbital elements from MPEC 1998-K11:

T = 1998 May 8.616 TT  ω = 110.^{o}563
Ω = 351.636 }2000.0
q = 0.15324 AU  i = 62.894
1998TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 21 5^{h}23.^{m}57 + 3^{o}16.^{′}9 0.843 0.482 28.^{o}3 95.^{o}8 4.5
23 5 37.14 − 0 41.8 0.849 0.539 32.1 90.9 5.0
25 5 49.76 − 4 26.1 0.861 0.594 35.8 86.2 5.4
27 6 01.59 − 7 55.0 0.877 0.647 39.2 81.8 5.8
29 6 12.78 −11 08.2 0.897 0.700 42.4 77.7 6.2
31 6 23.39 −14 06.3 0.920 0.750 45.3 73.9 6.6
June 2 6 33.52 −16 50.0 0.947 0.800 48.0 70.4 6.9
4 6 43.21 −19 20.3 0.975 0.848 50.4 67.2 7.2
6 6 52.52 −21 38.4 1.005 0.895 52.6 64.2 7.5

Visual m_{1} estimates (B = binoculars): May 17.35 UT, 2.8 (F. Farrell, Christies Beach, S. Australia, 7×50 B); 18.38, 3.5 (M. Mattiazzo, Adelaide, S. Australia, 7×50 B); 18.86, 3.7 (J. G. de S. Aguiar, Campinas, Brazil, 11×80 B); 19.68, 3.8 (T. Cooper, Bredell, Kempton Park, South Africa, 10×50 B); 20.91, 4.5 (N. Falsarella, São Jose do Rio Preto, Brazil, 7×50 B); 22.37, 4.5 (R. H. McNaught, Bugaldie, N.S.W., naked eye).

1998 May 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green